# MEPG: Multi-Expert Planning and Generation for Compositionally-Rich Image Generation

### v1

@@ 3.2 Mixture of Experts Diffusion ( MED ) Module
Through the weight allocation mechanism of the SparseMoeBlock, the gating function dynamically controls the participation of each expert model. We replace the qkv components of both the feed-forward ($F$) and attention (Attn) modules in denoising UNet with our SparseMoeBlock, where the experts are stacked from the $F$ or Attn modules of several base models, enabling model selection and weight distribution. Since our MED module does not modify the original model parameters, it supports the lightweight addition and removal of expert models.
@@ 3.3 Cross Denoising Scheduling
To balance global consistency and local detail control in image generation, we propose a cross denoising mechanism that combines global and local denoising strategies,which have been shown in figure 4 . Given a total of NN denoising steps, we divide the process into two phases: the first $p_{1}$ proportion of steps are dominated by local denoising, and the remaining $1-p_{1}$ proportion are dominated by global denoising.Specifically,let $t$ denote the current denoising step.The denoising stage at step $t$ is scheduled as follows:

### v2

@@ 3.2 Mixture of Experts Diffusion ( MED ) Module
Through the weight allocation mechanism of the SparseMoeBlock, the gating function dynamically controls the participation of each expert model. We replace the qkv components of both the feed-forward (FF) and attention (Attn) modules in denoising UNet with our SparseMoeBlock, where the experts are stacked from the FF or Attn modules of several base models, enabling model selection and weight distribution. Since our MED module does not modify the original model parameters, it supports the lightweight addition and removal of expert models.
@@ 3.3 Cross Denoising Scheduling
To balance global consistency and local detail control in image generation, we propose a cross denoising mechanism that combines global and local denoising strategies,which have been shown in figure 4 . Given a total of $N$ denoising steps, we divide the process into two phases: the first $p_{1}$ proportion of steps are dominated by local denoising, and the remaining $1-p_{1}$ proportion are dominated by global denoising.Specifically,let $t$ denote the current denoising step.The denoising stage at step $t$ is scheduled as follows: